OpenStudy (allgemeine-klause01):

The point-slope form of the equation of the line that passes through (–5, –1) and (10, –7) is y + 7 = (x – 10). What is the standard form of the equation for this line? any answer will be of help.

OpenStudy (anonymous):

First, do you know what standard form of a line looks like? As in just the general formula.

OpenStudy (aihberkhan):

Okay. So this is not really a calculation problem... it just wants to know the standard form of the equation.. So the standard form is: \[y - y1 = m(x - x1)\]

OpenStudy (aihberkhan):

Hope this helped! Have a great day and Happy Holidays! :) Also a medal would be much appreciated! Just click best response next to my answer. Thank You! Also, a fan will be much appreciated as well! Just hover over my avatar and click “Become a fan”. This will tell you whenever I am online so you can ask me for help in any of your other questions! If you see that I am online and need help with a question, just tag me in your question! @allgemeine-Klause01

OpenStudy (anonymous):

That is not the standard form of a line.

OpenStudy (allgemeine-klause01):

i know its just a simple standard form question but every time i do it myself it won't work out like the answers I'm given

OpenStudy (anonymous):

Alright so start with that, the standard form of a line is: \[Ax+By=C\] Where A, B, and C are constants. You currently have the line in point-slope form, which is: \[y-y _{1}=m(x-x _{1})\] What would you do to go from what you have to standard form?

OpenStudy (allgemeine-klause01):

make it into a slope intercept and from there into a standard form

OpenStudy (anonymous):

Alright, that's a great way of doing it, so multiply it out into slope-intercept form and then post it and I'll check it and we can go from there!

OpenStudy (allgemeine-klause01):

y-7=-1/3(x-6) == y=-1/3x+1 is what i got

OpenStudy (anonymous):

Just to double check, is this the same question as in the original post? You wrote originally: \[y+7=(x-10)\] And there's no use solving the wrong problem.

OpenStudy (allgemeine-klause01):

oops sorry hold on yeah sorry wrong problem

OpenStudy (anonymous):

No problem, I'm glad I caught it before we went further.

OpenStudy (allgemeine-klause01):

y=-2/5x-17 is what i got for the slope intercept

OpenStudy (anonymous):

There is an error somewhere in your work, how about you double check it and then if you can't find it I'll post my work to compare.

OpenStudy (allgemeine-klause01):

(srry the -2/5 didn't copy over something with code) y+7=-2/5(x-10) am i still off??

OpenStudy (allgemeine-klause01):

cause i have y=-2/5-17 i might be off i don't know

OpenStudy (anonymous):

Double check what you get on the right side when you distribute the -2/5

OpenStudy (allgemeine-klause01):

forgive me I've forgotten how to do that

OpenStudy (anonymous):

That's alright. I'll draw it out (forgive my drawing skills). First you multiply -2/5 times x, and then you multiply -2/5 times 10, and then put them together: